# Electrodynamics of Vortices in Quasi-2D Scalar Bose-Einstein Condensates

### v3

@@ I Introduction
Topological defects are ubiquitous in physics. A symmetry-breaking second-order phase transition generally leads to the formation of topological defects that can be classified according to the topology of the vacuum manifold using homotopy groups [1, 2, 3, 4, 5, 6, 7, 8]. This classification identifies different kinds of defects such as kinks, vortices, domain walls, skyrmions, etc. Among them, $U(1)$ vortices describe point-like singularities of a complex scalar field with quantized circulation. Their occurrence in Bose-Einstein condensates (BEC) indicates their superfluid character [9]. In such context, they can be created spontaneously by driving the transition from a normal fluid to a superfluid (e.g., by a thermal quench), as theoretically analyzed [10, 11] and experimentally demonstrated [12, 13, 14, 15]. An alternative mechanism to create vortices involves rotating the superfluid cloud, pumping angular momentum in the system [16, 17]. Vortices in degenerate ultracold gases can also be produced by phase imprinting [18, 19, 20], by merging independent BEC [21], by stirring laser beams [22, 23, 24, 25, 26], or by making superfluid flow pass an obstacle [27]. Progress in manipulating and controlling ultracold gases makes it possible to design arbitrary patterns of vortices in a BEC sample [28].
@@
Vortex-vortex interactions are known to scale logarithmically with the inter-vortex distance within some approximations. In the limit of a homogeneous BEC number density, excluding the core region of each vortex, an ensemble of static vortices in a non-rotating quasi-two-dimensional (quasi-2D) scalar BEC without dissipation can be mathematically regarded as a 2D Coulomb gas; see, e.g., Ref. [29].
@@
For non-rotating three-dimensional (3D) ${}^{4}\textrm{He}$ superfluid without dissipation, in the limit where the superfluid density is constant outside vortices, Refs. [30, 31] presented the duality between vortices in a thin cylindrical system and 2D electrodynamics. Refs. [32, 33] extended it to the general 3D case by using $\phi^{4}$ theory and minimizing the action with respect to the fluctuation of the superfluid density. In the absence of an external potential, Ref. [34] studied the asymptotic vortex dynamics in 2D nonlinear Schrödinger equation, and Ref. [35] constructed the duality between 2D electrodynamics and vortices in the (2+1) dimensional nonlinear wave equation. For the non-rotating scalar BEC without dissipation, when the BEC number density is approximately constant, it has been shown that the motion of the vortex can be described according to the nonrelativistic dynamics of stings in 3D system [36], and the effective Maxwell’s equations in a quasi-2D system [37]. This connection is valid for an inhomogeneous time-independent BEC number density in the case of a non-rotating quasi-2D scalar BEC [38].
@@
However, the BEC number density is zero at the core of a vortex, and thus, a vortex has a finite core size, which is about the order of the healing length (coherence length) $\xi_{h}\coloneqq\hbar/\sqrt{2Mgn_{m}}$, where $n_{m}$ is the mean BEC number density [39]. As a result, the fluctuation of the BEC number density cannot be neglected, especially around the core of the vortex. To simplify the problem, the point-vortex model (PVM) has been widely used. Yet, such description cannot account for the dynamic vortex creation and annihilation processes if one uses the mean-field Gross-Pitaevskii equation (GPE) under non-rotating quasi-2D scalar BEC without dissipation [40, 41, 42, 43]. It thus remains to be established whether the connection between vortices and effective Maxwell’s electrodynamics is valid in a quasi-2D scalar BEC with dissipation, beyond the PVM or under rotation.
@@
In this paper, starting from the microscopic Hermitian Hamiltonian of a non-rotating quasi-2D scalar bosons in the s-wave scattering limit [44, 45], we define the superfluid velocity via the probability current, and present the condition of the conservation of the topological charges of the vortices in general “beyond the GPE” case, i.e., when the mean-field limit of the Heisenberg equation of motion has a dissipative term (e.g., [46, 47, 48, 49, 50, 51, 52, 53]). Using that condition, we show how vortices in a quasi-2D scalar BEC can be connected to the effective Maxwell’s equations even beyond the PVM and without assuming a time-independent BEC number density. From such duality, we show that the phenomenological damped-PVM [54] can be induced and generalized beyond the PVM description and present how to calculate the temporal change of the circulation. Under the GPE + PVM, we show that the logarithmic vortex interaction may need correction if vortices move. We also show that one can recover previously known results when the fluctuations of the BEC number density are negligible.
@@ II Hamiltonian of quasi-2D scalar BEC and the continuity equation
We start by showing the relation between the mean-field limit of the Heisenberg equations of motion for non-rotating quasi-2D scalar BEC on the $xy$ plane in a region $\mathcal{A}$ and vortex quantization. We first review the case within the GPE description and then generalize it to the case with dissipation or rotation.
Let us introduce a unit vector $\bm{e}_{j}$ along the $+j$ axis ($j=x,y,z$), the position vector $\bm{r}\coloneqq\sum_{j=x,y}r_{j}\bm{e}_{j}$ on the $xy$ plane, and $\nabla\coloneqq\sum_{j=x,y}\bm{e}_{j}\partial/\partial r_{j}$. For later convenience, we define $\bm{e}_{\perp}\coloneqq\bm{e}_{z}$, $v\coloneqq\left|\bm{v}\right|=\sqrt{\bm{v}\cdot\bm{v}}$ for any vector $\bm{v}$, and $\left|\psi\right|\coloneqq\sqrt{\psi^{*}\psi}$ for any complex function $\psi$ with its complex conjugate being $\psi^{*}$. Definitions of symbols we used are summarized in Tables 1 and 2.
@@
BEC Bose-Einstein condensates.
GPE Gross-Pitaevskii equation.
PVM Point-vortex model.
$n$D $n$-dimensional.
$\mathcal{A}$ Region where the quasi-2D BEC is.
$M$ Mass of the boson.
$N$ Number of bosons in BEC.
$\hat{\psi}\left(\bm{r},t\right)$ Bosonic field operator.
$\hat{n}\left(\bm{r},t\right)$ Number density operator (see Eqs. (2)).
$\hat{\bm{J}}\left(\bm{r},t\right)$ Probability current operator (see Eqs. (2)).
$\hat{A}^{\dagger}$ Hermitian conjugate of the quantum operator $\hat{A}$.
$\hat{A}\pm h.c.$ $\hat{A}\pm\hat{A}^{\dagger}$, where $\hat{A}$ can be any quantum operator.
$A$ Quantum operator $\hat{A}$ in the mean-field limit.
$\psi^{*}$, $\left|\psi\right|$ Complex conjugate of $\psi$, and $\left|\psi\right|\coloneqq\sqrt{\psi^{*}\psi}$.
$A\pm c.c.$ $A\pm A^{*}$, where $A$ can be any complex function.
$v$ Magnitude of the real vector $\bm{v}$ ($v\coloneqq\sqrt{\bm{v}\cdot\bm{v}}$).
$\Phi_{K}\left(\bm{r},t\right)$ Mean-field kinetic energy density (See Eq. (8)).
$\Phi_{Q}\left(\bm{r},t\right)$ Quantum potential (See Eq. (9)).
$\bm{v}_{s}\left(\bm{r},t\right)$ Superfluid velocity (See Eq. (12)).
$\varphi\left(\bm{r},t\right)$ Phase of the mean-field wavefunction.
$\bm{f}_{K}\left(\bm{r},t\right)$ Effective kinetic force (See Eq. (18)).
$U_{\rm sf}\left(\bm{r},t\right)$ Effective potential (See Eq. (19)).
@@
(Dissipation, rotation, etc. See section II.2).
$\bm{e}_{\perp}$ Unit vector perpendicular to the system in $\mathcal{A}$.
$q_{j}$ Topological charge of the vortex.
$\mathbb{Z}$ Set of integers.
$\bm{r}_{\alpha_{j}}$ Position of the core of the vortex (See Eq. (30)).
@@
(See Eq. (30)).
$\partial\mathcal{M}$ Boundary of the region $\mathcal{M}$.
@@
$\delta\left(\bm{r}\right)$ Dirac delta function.
$\left|\mathcal{M}\right|$ Area of the region $\mathcal{M}$.
$\bar{U}_{\rm sf}\left(t\right)$ Spatial average of $U_{\rm sf}\left(\bm{r},t\right)$ (See Eq. (45)).
$\theta\left(x\right)$ Heaviside step function.
$\mathcal{D}^{D}_{R}\left(\bm{r}\right)$ $D$-dimensional disk (radius $R$) centered at $\bm{r}$.
@@
$\rho_{v}\left(\bm{r},t\right)$ Vortex charge density (See Eq. (35)).
@@
(See Eqs. (37) and (52) for its meaning).
$\epsilon_{\rm sf}$ Effective vacuum permittivity.
@@
(See Eqs. (38)).
$\bm{P}_{\rm sf}\left(\bm{r},t\right)$ Effective polarization density (See Eqs. (38)).
@@
(Maximum speed of sound in scalar BEC).
$\bm{S}_{\rm sf}\left(\bm{r},t\right)$ Effective Poynting vector (See Eq. (52)).
$V_{e,\rm{sf}}\left(\bm{r},t\right)$ Effective electric potential (See Eqs. (53)).
@@
(See Eq. (LABEL:def_Bsf) and section V).
II.1 Case 1: Gross-Pitaevskii equation for non-rotating scalar BEC
Using second quantization and the $s$-wave scattering limit, the Hamiltonian $\hat{H}\left(t\right)$ in the Heisenberg picture of a non-rotating quasi-2D scalar Bose gas on the $xy$ plane in a region $\mathcal{A}$ can be expressed as [44, 45]
@@ II.1 Case 1: Gross-Pitaevskii equation for non-rotating scalar BEC
$\frac{\hat{\bm{J}}\left(\bm{r},t\right)}{\partial t}$ $=$ $)\displaystyle\frac{\hbar^{2}}{4M^{2}}\left\lgroup\left\{\nabla\left[\nabla^{2}\hat{\psi}^{\dagger}\left(\bm{r},t\right)\right]\right\}\hat{\psi}\left(\bm{r},t\right)+h.c.\right\rgroup$ (4)
@@
To simplify the problem, we will focus on the mean-field limit where the BEC order parameter is given by $\psi\left(\bm{r},t\right)\coloneqq\left\langle\hat{\psi}\left(\bm{r},t\right)\right\rangle$, satisfying $\psi^{\dagger}\left(\bm{r},t\right)\psi\left(\bm{r},t\right)=\psi\left(\bm{r},t\right)\psi^{\dagger}\left(\bm{r},t\right)$ as we consider scalar (single-component) BEC. In this mean-field limit, one can introduce the BEC number density $n\left(\bm{r},t\right)\coloneqq\left|\psi\left(\bm{r},t\right)\right|^{2}$ and the mean-field probability current
@@
which takes the same form as the Heisenberg equation of motion for $\hat{\psi}\left(\bm{r},t\right)$ upon the replacement $\hat{\psi}\left(\bm{r},t\right)\rightarrow\psi\left(\bm{r},t\right)$. From the Hamiltonian in Eq. (1), we define the mean-field energy $E\left(t\right)$ of the non-rotating quasi-2D scalar BEC as
@@
wherever the superfluid density is finite $n\left(\bm{r},t\right)\neq 0$, the vorticity vanishes, ×$\nabla\times\bm{v}_{s}\left(\bm{r},t\right)=0$. Conversely, the curl of $\bm{v}_{s}\left(\bm{r},t\right)$ need not vanish when $n\left(\bm{r},t\right)=0$. In section III, we will determine the curl of $\bm{v}_{s}\left(\bm{r},t\right)$ at $n\left(\bm{r},t\right)=0$ by using the vortex quantization.
@@
For finite density $n\left(\bm{r},t\right)\neq 0$, one can see that $\bm{f}_{K}\left(\bm{r},t\right)=-\nabla\left[\Phi_{Q}\left(\bm{r},t\right)+Mv^{2}_{s}\left(\bm{r},t\right)/2\right]$. However, the effective kinetic force $\bm{f}_{K}\left(\bm{r},t\right)$ is singular at $n\left(\bm{r},t\right)=0$. In section III, we will show how to determine the curl of $\bm{f}_{K}\left(\bm{r},t\right)$, leading to ×$\nabla\times\bm{f}_{K}\left(\bm{r},t\right)\neq 0$ when vortices are moving.
@@
II.2 Case 2: beyond the Gross-Pitaevskii equation
The Gross-Pitaevskii equation in Eq. (6) is obtained by neglecting the bosonic field operator of the noncondensed particles [45]. At nonzero temperature, a dissipation term emerges (e.g., [46, 47, 48, 49, 50, 51, 52, 53]). Let us generalize Eqs. (10) and (11) as

### v4

# Electrodynamics of vortices in quasi-two-dimensional scalar Bose-Einstein condensates

@@ I Introduction
Topological defects are ubiquitous in physics. A symmetry-breaking second-order phase transition generally leads to the formation of topological defects that can be classified according to the topology of the vacuum manifold using homotopy groups [1, 2, 3, 4, 5, 6, 7, 8]. This classification identifies different kinds of defects such as kinks, vortices, domain walls, skyrmions, etc. Among them, $U(1)$ vortices describe pointlike singularities of a complex scalar field with quantized circulation [9]. Their occurrence in Bose-Einstein condensates (BEC) indicates their superfluid character [10, 11]. In such context, they can be created spontaneously by driving the transition from a normal fluid to a superfluid (e.g., by a thermal quench), as theoretically analyzed [12, 13] and experimentally demonstrated [14, 15, 16, 17]. An alternative mechanism to create vortices involves rotating the normal cloud [18] or the superfluid cloud, pumping angular momentum in the system [19, 20]. Vortices in degenerate ultracold gases can also be produced by phase imprinting [21, 22, 23], by merging independent BEC [24], by stirring laser beams [25, 26, 27, 28, 29], or by making superfluid flow pass an obstacle [30]. Progress in manipulating and controlling ultracold gases makes it possible to design arbitrary patterns of vortices in a BEC sample [31].
@@
Vortex-vortex interactions are known to scale logarithmically with the inter-vortex distance within some approximations. In the limit of a homogeneous BEC number density, excluding the core region of each vortex, an ensemble of static vortices in a nonrotating quasi-two-dimensional (quasi-2D) scalar BEC without dissipation can be mathematically regarded as a 2D Coulomb gas [9].
@@
For nonrotating three-dimensional (3D) ${}^{4}\textrm{He}$ superfluid without dissipation, a duality between vortices in a thin cylindrical system and 2D electrodynamics has been reported in the limit where the superfluid density is constant outside vortices [32, 33]. Such a duality has been extended to the general 3D case by using $\phi^{4}$ theory and minimizing the action with respect to the fluctuation of the superfluid density [34, 35]. In the absence of an external potential, the asymptotic vortex dynamics in 2D nonlinear Schrödinger equation has been studied [36], and a duality between 2D electrodynamics and vortices in the (2+1) dimensional nonlinear wave equation has been put forward [37]. For the nonrotating scalar BEC without dissipation, when the BEC number density is approximately constant, it has been shown that the motion of the vortex can be described according to the nonrelativistic dynamics of strings in 3D system [38], and the effective Maxwell’s equations in a quasi-2D system [39]. This connection is valid for an inhomogeneous time-independent BEC number density in the case of a nonrotating quasi-2D scalar BEC [40].
@@
However, the BEC number density is zero at the core of a vortex, and thus, a vortex has a finite core size, which is about the order of the healing length (coherence length) $\xi_{h}\coloneqq\hbar/\sqrt{2Mgn_{m}}$, where $n_{m}$ is the mean BEC number density [41]. As a result, the fluctuation of the BEC number density cannot be neglected, especially around the core of the vortex. To simplify the problem, the point-vortex model (PVM) has been widely used. Yet, such description cannot account for the dynamic vortex creation and annihilation processes if one uses the mean-field Gross-Pitaevskii equation (GPE) under nonrotating quasi-2D scalar BEC without dissipation [42, 43, 44, 45]. It thus remains to be established whether the connection between vortices and effective Maxwell’s electrodynamics is valid in a quasi-2D scalar BEC with dissipation, beyond the PVM or under rotation.
@@
In this paper, starting from the microscopic Hermitian Hamiltonian of a nonrotating quasi-2D scalar bosons in the s-wave scattering limit [11, 46], we define the superfluid velocity via the probability current, and present the condition of the conservation of the topological charges of the vortices in general “beyond the GPE” case, i.e., when the mean-field limit of the Heisenberg equation of motion has a dissipative term (e.g., Refs. [47, 48, 49, 50, 51, 52, 53, 54]). Using that condition, we show how vortices in a quasi-2D scalar BEC can be connected to the effective Maxwell’s equations even beyond the PVM and without assuming a time-independent BEC number density. From such duality, we show that the damped-PVM [55, 56] can be alternatively induced and generalized beyond the PVM description and present how to calculate the temporal change of the circulation. Under the GPE + PVM, we show that the logarithmic vortex interaction may need correction if vortices move. We also show that one can recover previously known results when the fluctuations of the BEC number density are negligible. For a quick reference, we present the key ideas on the duality we constructed in Fig. 1.
Figure 1: Duality between the vortices in the quasi-2D scalar BEC and the electrodynamics in the matter. The derivations are in Sec. IV, and the full effective Maxwell’s equations are shown in Table 3. On the right-hand side, $\bm{D}\left(\bm{r},t\right)$ is the electric displacement field, $\bm{H}\left(\bm{r},t\right)$ is the magnetic field strength, $Q_{j}\left(t\right)$ is the free electric charge with index $j$ at time $t$, and $\bm{J}_{f}\left(\bm{r},t\right)$ is the free electric current density. For other definitions of symbols, refer to Tables 1 and 2.
@@ II Hamiltonian of quasi-2D scalar BEC and the continuity equation
We start by showing the relation between the mean-field limit of the Heisenberg equations of motion for nonrotating quasi-2D scalar BEC on the $xy$ plane in a region $\mathcal{A}$ and vortex quantization. We first review the case within the GPE description and then generalize it to the case with dissipation or rotation.
Let us introduce a unit vector $\bm{e}_{j}$ along the $+j$ axis ($j=x,y,z$), the position vector $\bm{r}\coloneqq\sum_{j=x,y}r_{j}\bm{e}_{j}$ on the $xy$ plane, and $\nabla\coloneqq\sum_{j=x,y}\bm{e}_{j}\partial/\partial r_{j}$. For later convenience, we define $\bm{e}_{\perp}\coloneqq\bm{e}_{z}$, $v\coloneqq\left|\bm{v}\right|=\sqrt{\bm{v}\cdot\bm{v}}$ for any vector $\bm{v}$, and $\left|\psi\right|\coloneqq\sqrt{\psi^{*}\psi}$ for any complex function $\psi$ with its complex conjugate being $\psi^{*}$. For convenience, a summary of the symbols we used is provided in Tables 1 and 2.
@@
BEC Bose-Einstein condensates
GPE Gross-Pitaevskii equation
PVM Point-vortex model
$n$D $n$-dimensional
$\mathcal{A}$ Region where the quasi-2D BEC is
$M$ Mass of the boson
$N$ Number of bosons in BEC
$\hat{\psi}\left(\bm{r},t\right)$ Bosonic field operator
$\hat{n}\left(\bm{r},t\right)$ Number density operator [see Eqs. (2)]
$\hat{\bm{J}}\left(\bm{r},t\right)$ Probability current operator [see Eqs. (2)]
$\hat{A}^{\dagger}$ Hermitian conjugate of the quantum operator $\hat{A}$
$\hat{A}\pm h.c.$ $\hat{A}\pm\hat{A}^{\dagger}$, where $\hat{A}$ can be any quantum operator
$A$ Quantum operator $\hat{A}$ in the mean-field limit
$\psi^{*}$, $\left|\psi\right|$ Complex conjugate of $\psi$, and $\left|\psi\right|\coloneqq\sqrt{\psi^{*}\psi}$
$A\pm c.c.$ $A\pm A^{*}$, where $A$ can be any complex function
$v$ Magnitude of the real vector $\bm{v}$ ($v\coloneqq\sqrt{\bm{v}\cdot\bm{v}}$)
$\Phi_{K}\left(\bm{r},t\right)$ Mean-field kinetic energy density [See Eq. (8)]
$\Phi_{Q}\left(\bm{r},t\right)$ Quantum potential [See Eq. (9)]
$\bm{v}_{s}\left(\bm{r},t\right)$ Superfluid velocity [See Eq. (12)]
$\varphi\left(\bm{r},t\right)$ Phase of the mean-field wavefunction
$\bm{f}_{K}\left(\bm{r},t\right)$ Effective kinetic force [See Eq. (18)]
$U_{\rm sf}\left(\bm{r},t\right)$ Effective potential [See Eq. (19)]
@@
(Dissipation, rotation, etc. See Sec. II.2)
$\bm{e}_{\perp}$ Unit vector perpendicular to the system in $\mathcal{A}$
$q_{j}$ Topological charge of the vortex
$\mathbb{Z}$ Set of integers
$\bm{r}_{\alpha_{j}}$ Position of the core of the vortex [See Eq. (30)]
@@
[See Eq. (30)]
$\partial\mathcal{M}$ Boundary of the region $\mathcal{M}$
@@
$\delta\left(\bm{r}\right)$ Dirac delta function
$\left|\mathcal{M}\right|$ Area of the region $\mathcal{M}$
$\bar{U}_{\rm sf}\left(t\right)$ Spatial average of $U_{\rm sf}\left(\bm{r},t\right)$ [See Eq. (45)]
$\theta\left(x\right)$ Heaviside step function
$\mathcal{D}^{D}_{R}\left(\bm{r}\right)$ $D$-dimensional disk (radius $R$) centered at $\bm{r}$
@@
$\rho_{v}\left(\bm{r},t\right)$ Vortex charge density [See Eq. (35)]
@@
[See Eqs. (37) and (52) for its meaning]
$\epsilon_{\rm sf}$ Effective vacuum permittivity
@@
[See Eqs. (38)]
$\bm{P}_{\rm sf}\left(\bm{r},t\right)$ Effective polarization density [See Eqs. (38)]
@@
(Maximum speed of sound in scalar BEC)
$\bm{S}_{\rm sf}\left(\bm{r},t\right)$ Effective Poynting vector [See Eq. (52)]
$V_{e,\rm{sf}}\left(\bm{r},t\right)$ Effective electric potential [See Eqs. (53)]
@@
[See Eq. (LABEL:def_Bsf) and Sec. V]
II.1 Case 1: Gross-Pitaevskii equation for nonrotating scalar BEC
Using second quantization and the $s$-wave scattering limit, the Hamiltonian $\hat{H}\left(t\right)$ in the Heisenberg picture of a nonrotating quasi-2D scalar Bose gas on the $xy$ plane in a region $\mathcal{A}$ can be expressed as [11, 46]
@@ II.1 Case 1: Gross-Pitaevskii equation for nonrotating scalar BEC
$\frac{\partial\hat{\bm{J}}\left(\bm{r},t\right)}{\partial t}$ $=$ $)\displaystyle\frac{\hbar^{2}}{4M^{2}}\left\lgroup\left\{\nabla\left[\nabla^{2}\hat{\psi}^{\dagger}\left(\bm{r},t\right)\right]\right\}\hat{\psi}\left(\bm{r},t\right)+h.c.\right\rgroup$ (4)
@@
To simplify the problem, we focus on the mean-field limit where the BEC order parameter is given by $\psi\left(\bm{r},t\right)\coloneqq\left\langle\hat{\psi}\left(\bm{r},t\right)\right\rangle$, satisfying $\psi^{\dagger}\left(\bm{r},t\right)\psi\left(\bm{r},t\right)=\psi\left(\bm{r},t\right)\psi^{\dagger}\left(\bm{r},t\right)$ as we consider a scalar (single-component) BEC. In this mean-field limit, one can introduce the BEC number density $n\left(\bm{r},t\right)\coloneqq\left|\psi\left(\bm{r},t\right)\right|^{2}$ and the mean-field probability current
@@
which takes the same form as the Heisenberg equation of motion for $\hat{\psi}\left(\bm{r},t\right)$ upon the replacement $\hat{\psi}\left(\bm{r},t\right)\rightarrow\psi\left(\bm{r},t\right)$. From the Hamiltonian in Eq. (1), we define the mean-field energy $E\left(t\right)$ of the nonrotating quasi-2D scalar BEC as
@@
wherever the superfluid density is finite $n\left(\bm{r},t\right)\neq 0$, the vorticity vanishes, ×$\nabla\times\bm{v}_{s}\left(\bm{r},t\right)=0$. Conversely, the curl of $\bm{v}_{s}\left(\bm{r},t\right)$ need not vanish when $n\left(\bm{r},t\right)=0$. In Sec. III, we will determine the curl of $\bm{v}_{s}\left(\bm{r},t\right)$ at $n\left(\bm{r},t\right)=0$ by using the vortex quantization.
@@
For finite density $n\left(\bm{r},t\right)\neq 0$, one can see that $\bm{f}_{K}\left(\bm{r},t\right)=-\nabla\left[\Phi_{Q}\left(\bm{r},t\right)+Mv^{2}_{s}\left(\bm{r},t\right)/2\right]$. However, the effective kinetic force $\bm{f}_{K}\left(\bm{r},t\right)$ is singular at $n\left(\bm{r},t\right)=0$. In Sec. III, we will show how to determine the curl of $\bm{f}_{K}\left(\bm{r},t\right)$, leading to ×$\nabla\times\bm{f}_{K}\left(\bm{r},t\right)\neq 0$ when vortices are moving.
@@
II.2 Case 2: beyond the nonrotating Gross-Pitaevskii equation
The Gross-Pitaevskii equation in Eq. (6) is obtained by neglecting the bosonic field operator of the noncondensed particles [46] in the nonrotating case. However, at nonzero temperature, a dissipation term emerges (e.g., [47, 48, 49, 50, 51, 52, 53, 54]). To go beyond the nonrotating dissipationless case, let us generalize Eqs. (10) and (11) as